# Text-to-SQL Task-oriented Dialogue Ontology Construction

### v1

@@ Abstract
We introduce TeQoDO: a Text-to-SQL task-oriented Dialogue Ontology construction method. Here, an LLM autonomously builds a TOD ontology from scratch without supervision using its inherent SQL programming capabilities combined with dialogue theory provided in the prompt.
We show that TeQoDO outperforms transfer learning approaches, and its constructed ontology is competitive on a downstream dialogue state tracking task. Ablation studies demonstrate the key role of dialogue theory. TeQoDO also scales to allow construction of much larger ontologies, which we investigate on a Wikipedia and ArXiv dataset. We view this as a step towards broader application of ontologies to increase LLM explainability. [^1] Code will be released upon publication.
@@ 1 Introduction
Large language models (LLMs) have become ubiquitous knowledge processing systems, with the ability to reach or surpass human performance on a wide variety of natural language processing tasks. These models are pre-trained on massive corpora and aligned via human feedback using reinforcement learning Brown et al. (2020); Ouyang et al. (2022). Despite these remarkable abilities, there are some inherent problems associated with these systems. Their knowledge is stored in vast numbers of parameters, which makes it very difficult to understand their behaviour, even via probing Cífka and Liutkus (2023). It is therefore extremely challenging to verify their inherent knowledge Zhong et al. (2024). Moreover, LLMs often produce confident but non-factual outputs that just appear plausible on a superficial level Sahoo et al. (2024); Feng et al. (2024). Finally, many LLMs and their training data are closed-source.
@@
Ontologies provide a human-readable means of reconstructing knowledge-based reasoning in language models Gruber (1995); Lo et al. (2024). Manually building ontologies for all relevant domains is infeasible Milward and Beveridge (2003). To ensure broad applicability, ontologies should be constructed automatically, reliably, and consistently. Automatic ontology construction presents a promising solution to these challenges.
@@
In this work, we focus on constructing task-oriented dialogue (TOD) ontologies from raw dialogue data. The goal is to extract relevant information and organise it into a meaningful hierarchy for handling user queries. TOD ontologies consist of domains, slots, and values, with system actions and user intents defined over this domain-slot-value structure. Figures˜1 and 2 illustrate ontology construction in this setting. Such ontologies are essential for modern TOD systems Young et al. (2013); Hudeček and Dusek (2023).
@@
Most TOD ontology construction approaches follow two steps Hudeček et al. (2021); Vukovic et al. (2022): (1) term extraction from dialogue data and (2) relation extraction between the terms from the previous step. Existing approaches tackle these steps separately and rely on annotated training data Vukovic et al. (2024); Finch et al. (2024). This has the downside of additional and potentially error-prone training, and the danger of information loss between the two processing steps.
@@
We propose TeQoDO, a Text-to-SQL task-oriented Dialogue Ontology construction approach. Figure˜3 provides an overview of the method. TeQoDO uses LLMs’ code understanding and generation capabilities Chen et al. (2023) to build ontologies from scratch via SQL. It incrementally constructs the database by iterating over dialogues, retrieving relevant existing content, and updating the DB with new information. The model uses state tracking from dialogue theory to distinguish new content from existing data in the DB. The update prompt includes a notion of success to help align schema changes with the user’s goals. Query results are augmented with semantically similar concepts or example values from existing columns to improve coverage and consistency.
@@
Text-to-SQL provides a structured format familiar to LLMs through code encountered during pre-training Deng et al. (2022); Zhao et al. (2024). It reduces the need for textual prompts, as SQL tables, columns, and values align with TOD ontology domains, slots, and values. To our knowledge, we are the first to use text-to-SQL to build a TOD ontology from scratch. We also adapt similarity-based evaluation metrics for ontology learning Lo et al. (2024) to reflect the hierarchical nature of TOD ontologies. This enables evaluation of various concept types, such as domains and slots.
@@
We run experiments on two widely used TOD datasets: MultiWOZ Eric et al. (2020) and Schema-Guided Dialogue (SGD) Rastogi et al. (2020). TeQoDO outperforms recent TOD ontology construction methods Finch et al. (2024); Vukovic et al. (2024). To test generalisation, we apply TeQoDO to general ontology datasets from Lo et al. (2024) and show it scales to large ontologies. In summary, our contributions are:
@@ 2.1 Ontology Construction
Figure 1: An example of ontology construction shows the extraction of the domain “Hotels”, with the value “cheap” assigned to the slot “Price”. System actions and user intents are defined based on the domains and slots specified in the ontology.
Prior to deep learning, ontology construction relied on frequency-based rules and linguistic features. With language models, clustering and supervised methods enabled feature-based learning, as outlined below.
@@
ontology construction (Frantzi and Ananiadou, 1999; Nakagawa and Mori, 2002) relies on frequency-based methods to extract key subsequences from text. Wermter and Hahn (2006) use linguistic features for more advanced term extraction. While interpretable, these methods are hard to adapt across domains and often yield low precision.
@@
approaches, such as Yu et al. (2022), apply unsupervised parsing and hierarchical clustering to group extracted terms into domains and slots. Finch et al. (2024) train a generative slot induction model to extract domain-slot-value triples, which are then clustered into a slot hierarchy. These methods usually produce many slots, whose interpretation depends entirely on how well they match the ground truth. This is because evaluation compares slots solely based on their values. They also rely heavily on data quality, embedding representations, and sensitive hyperparameters, such as the minimum cluster size.
@@
Lo et al. (2024) fine-tune LLMs to predict ontology sub-graphs at the document level and merge them using frequency-based rules to construct ontologies for Wikipedia and ArXiv datasets. They also propose evaluation functions covering both verbatim content and higher-level structural semantics compared to the ground truth, which we adapt to TOD. Vukovic et al. (2024) train a model to predict TOD ontology subgraphs at the dialogue level and aggregate them by merging all predictions. They improve generalisability by updating the decoding with constraints and confidence-based adjustments. However, their method focuses only on relation extraction, using ground truth terms as input. These methods require costly annotated training data from Budzianowski et al. (2018).
In contrast, our approach requires no training and conducts both term and relation extraction. We use Vukovic et al. (2024) and Finch et al. (2024) as competitive baselines.
@@ 2.2 Text-to-SQL
Li et al. (2024) present a large benchmark for text-to-SQL queries on big databases and find LLMs underperform compared to humans. Zhou et al. (2024) apply LLMs to database tasks such as query rewriting and index tuning using automatic prompt generation, DB-specific pre-training, model design, and fine-tuning. Yang et al. (2024) enhance text-to-SQL parsing for smaller models by synthesising training data with larger models. Allemang and Sequeda (2024) improve text-to-SPARQL Polleres (2014) by using a rule-based query checker leveraging the underlying ontology, with an LLM repairing queries based on this feedback. In contrast to our work, these methods work with existing ontologies and do not build ontologies from scratch using text-to-SQL.
@@ 3.1 Problem Formulation
Given a task-oriented dialogue dataset $D=\{d_{1},\ldots,d_{k}\}$, each dialogue $d_{i}$ is comprised of alternating user and system turns $\{u_{1},s_{1},\ldots,u_{m_{i}},s_{m_{i}}\}$ containing unstructured information about user-queried entities. The goal is to induce an ontology $O_{D}$ for the dataset. The ontology is a directed graph $O_{D}=(V,E)$ with five node types $V=V_{\text{domain}}\cup V_{\text{user\_intent}}\cup V_{\text{system\_action}}\cup V_{\text{slot}}\cup V_{\text{value}}$, forming a three-level hierarchy. The edge set $E$ is a subset of edges between specific node types, $E\subseteq(V_{\mathrm{domain}}\times V_{\mathrm{slot}})\cup(V_{\mathrm{slot}}\times V_{\mathrm{value}})\cup(V_{\mathrm{user\_intent}}\times V_{\mathrm{user\_intent}})\cup(V_{\mathrm{system\_action}}\times V_{\mathrm{system\_action}}).$
@@ 3.2 Method
Our main prompting approach combines task-oriented dialogue modelling with SQL, as outlined below. We then detail the steps of our iterative ontology construction pipeline. We utilise an LLM that creates SQL queries, which are then executed using Python.
@@
SQL is a query language for relational databases Chamberlin and Boyce (1974). We use SQLite, [^2] https://www.sqlite.org a serverless relational DB management system suitable for small-scale databases and schema-based knowledge representation. SQLite stores data in tables with columns and values, e.g., a “restaurant” table with a “price” column and “expensive” as a value. This structure aligns with TOD ontology formats, making it well-suited for ontology construction. For intents and actions, there are separate tables that store names of intent and action entries.
@@
We incorporate two concepts from modular task-oriented dialogue models (Young et al., 2013) into our prompt to improve the quality of generated update queries. First, dialogue state tracking (DST) is used as a separate step to distinguish existing database content from new input, based on the current schema. Second, we prompt the model to consider dialogue success, guiding it to make updates that support achieving the user goal. For a more formal description, see Table˜7 in the appendix.
@@
See Figure˜3 for an overview of TeQoDO with example queries and results. More detailed pipeline steps are given in Algorithm˜1 and are described in the following paragraphs.
@@
To ensure structural and naming consistency in TeQoDO, the model is first prompted to query the current DB tables. Initially, as the ontology is built from scratch, the DB is empty, and thus no results are returned. The resulting table list $T=\{t_{1},\ldots,t_{n}\}$ is included with the dialogue in step 2. The model is prompted to query column info for tables relevant to the dialogue. We then append the column query results to the prompt so the model generates SELECT queries consistent with the DB state. In step 3, it generates SELECT queries to retrieve specific entity values from the dialogue.
@@
In this step, the model restructures DB query results into dialogue state tracking style labels. It predicts the current dialogue’s value for each table-column pair and compares it with the DB’s stored information. This improves the distinction between existing DB information and new information from the dialogue.
The model is prompted to summarise dialogue states using the DB’s table and column names, as well as intents and actions. This summary includes information from both the user and the system. By predicting the state, the model condenses the dialogue’s information into a concise, model-determined format not fully detailed in the prompt. Notably, the model tracks the entire dialogue at once, not turn-by-turn. It must track all mentioned information, not only what the user queries.
Step 3: DB Update
In the final step, the model generates database update queries based on the current DB state and dialogue information. It is prompted to create update queries considering DB query and DST results. Update queries include CREATE TABLE, ALTER TABLE to add columns, and UPDATE to add values. The model must follow the existing DB structure and generate consistent updates, preferring existing tables over creating new ones. Additionally, the model is prompted to update the DB to support fulfilling the user’s goal by incorporating dialogue success into the prompt.
@@ 4.1 Datasets
The first dataset we employ is MultiWOZ 2.1 (Eric et al., 2020), a large-scale, multi-domain dialogue dataset containing human-human conversations annotated with domain, slot, and value labels. It covers information such as hotel bookings, restaurant reservations, taxi services, and attractions. We utilise the test set with 1,000 dialogues and 6 domains.
@@ 7  Generate update queries $U_{d_{i}}$ for success using $p=p+ST_{d_{i}}$
Second, we use the schema-guided dialogue (SGD) dataset (Rastogi et al., 2020). It is a diverse, multi-domain dataset for TOD modelling, featuring detailed schema annotations. It covers services like flight booking, calendar scheduling, banking, and media services, including unseen services at test time. We use the SGD test split for evaluation in the main results, containing 4,201 dialogues and 18 domains. We load these with ConvLab-3 Zhu et al. (2023).
@@
To test whether TeQoDO’s concepts apply beyond TOD data, we apply it to the Wikipedia and ArXiv datasets by Lo et al. (2024). The Wikipedia test set contains 242,148 article titles and abstracts; its gold ontology graph has 4 hierarchy levels with 8,033 nodes and 14,673 edges. Topics include various types of “injuries” or “legislative bodies”.
The ArXiv test set has 27,630 title-abstract pairs, 2 hierarchy levels with 61 nodes and 61 edges. Main topics are ArXiv categories like “Mathematics” with subcategories such as “Commutative Algebra”, making this ontology more abstract and higher level than Wikipedia’s. These ontologies do not follow the domain-slot-value hierarchy resembling SQL’s table-column-value structure. TeQoDO is applied to the test sets of both datasets.

### v2

@@ Abstract
We introduce TeQoDO: a Text-to-SQL task-oriented Dialogue Ontology construction method. Here, an LLM autonomously builds a TOD ontology from scratch using only its inherent SQL programming capabilities combined with concepts from modular TOD systems provided in the prompt.
We show that TeQoDO outperforms transfer learning approaches, and its constructed ontology is competitive on a downstream dialogue state tracking task. Ablation studies demonstrate the key role of modular TOD system concepts. TeQoDO also scales to allow construction of much larger ontologies, which we investigate on a Wikipedia and arXiv dataset. We view this as a step towards broader application of ontologies.  [^1] Code is available under https://gitlab.cs.uni-duesseldorf.de/general/dsml/teqodo-code-public
@@ 1 Introduction
Large language models (LLMs) have become ubiquitous knowledge processing systems, with the ability to reach or surpass human performance on a wide variety of natural language processing tasks. These models are pre-trained on massive corpora and aligned via human feedback using reinforcement learning brown2020gpt3; ouyang2022instructgpt. Despite these remarkable abilities, there are some inherent problems associated with these systems. Their knowledge is stored in vast numbers of parameters, which makes it very difficult to understand their behaviour, even via probing cifka-liutkus-2023-black. It is therefore extremely challenging to verify their inherent knowledge zhong2024knowledge. Moreover, LLMs often produce plausible-looking hallucinations sahoo-etal-2024-comprehensive; feng-etal-2024-infusing. Finally, many LLMs and their training data are closed-source.
@@
Ontologies provide a human-readable means of reconstructing knowledge-based reasoning in language models (LMs) by capturing knowledge in a structured and explicit manner gruber1995ontologydesign; lo2024ontologylearning. Manually building ontologies for all relevant domains is infeasible milward2003ontology. To ensure broad applicability, ontologies should be constructed automatically and consistently. Automatic ontology construction is a promising solution to these challenges.
@@
In this work, we focus on constructing task-oriented dialogue (TOD) ontologies from dialogues. The goal is to extract information and organise it into a hierarchy for handling user queries. TOD ontologies consist of domains, slots, and values, with system actions and user intents. Figure˜1 illustrates ontology construction in this setting. Ontologies are essential for TOD systems young2013pomdp; hudecek-dusek-2023-large.
@@
Most TOD ontology construction approaches follow two steps hudecek-etal-2021-discovering; vukovic-etal-2022-dialogue: (1) term extraction from dialogue data and (2) relation extraction between the terms from the previous step. Existing approaches tackle these steps separately and rely on annotated training data finch-etal-2024-transforming; vukovic-etal-2024-dialogue. This has the downside of additional and potentially error-prone training, and the danger of information loss between the two processing steps.
@@
We propose TeQoDO, a Text-to-SQL task-oriented Dialogue Ontology construction approach. Figure˜3 provides an overview of the method. TeQoDO uses LLMs’ code understanding and generation capabilities chen2022program to build ontologies from scratch via SQL. It incrementally constructs the database by iterating over dialogues, retrieving relevant existing content, and updating the DB with new information. The model uses state tracking from modular TOD to distinguish new content from existing data in the DB. The update prompt includes a notion of success to help align schema changes with the user’s goals. Query results are augmented with semantically similar concepts or example values from existing columns to improve coverage and consistency.
@@
Text-to-SQL provides a structured format familiar to LLMs through code encountered during pre-training deng-etal-2022-recent; zhao-etal-2024-deciphering. It reduces the need for textual prompts, as SQL tables, columns, and values align with TOD ontology domains, slots, and values. To our knowledge, we are the first to use text-to-SQL to build a TOD ontology from scratch. We also adapt similarity-based evaluation metrics for ontology learning lo2024ontologylearning to reflect the hierarchical nature of TOD ontologies. This enables evaluation of various concept types, such as domains and slots.
@@
We run experiments on two widely used TOD datasets: MultiWOZ eric-etal-2020-multiwoz and Schema-Guided Dialogue (SGD) rastogi2020towards. TeQoDO outperforms recent TOD ontology construction methods finch-etal-2024-transforming; vukovic-etal-2024-dialogue. To test generalisation, we apply TeQoDO to general ontology datasets from lo2024ontologylearning and show it scales to large ontologies. In summary, our contributions are:
@@ 2.1 Ontology Construction
Figure 1: Example ontology construction shows the extraction of the domain “Hotels”, with slot “Price” and value “cheap”. Actions and intents are defined based on the domains and slots.
Traditionally, ontologies are handcrafted. Prior to deep learning, efforts for automatic ontology construction relied on frequency-based rules and linguistic features. With LMs, clustering and supervised methods enabled feature-based learning.
@@
ontology construction (frantzi1999c; nakagawa2002simple) relies on frequency-based methods to extract key subsequences from text. wermter2006you use linguistic features for more advanced term extraction. While interpretable, these methods are hard to adapt across domains and often yield low precision.
@@
approaches, such as yu-etal-2022-unsupervised, apply unsupervised parsing and hierarchical clustering to group extracted terms into domains and slots. finch-etal-2024-transforming train a generative slot induction model to extract domain-slot-value triples, which are then clustered into a slot hierarchy. These methods usually produce many slots, whose interpretation depends entirely on how well they match the ground truth. They also rely heavily on data quality, embedding representations, and sensitive hyperparameters, such as the minimum cluster size.
@@
lo2024ontologylearning fine-tune LLMs to predict ontology sub-graphs at the document level and merge them using frequency-based rules to construct ontologies for Wikipedia and arXiv datasets. They also propose evaluation functions covering both verbatim content and higher-level structural semantics compared to the ground truth, which we adapt to TOD. vukovic-etal-2024-dialogue train a model to predict TOD ontology subgraphs at the dialogue level and aggregate them by merging all predictions. They improve generalisability by updating the decoding with constraints and confidence-based adjustments. However, their method focuses only on relation extraction, using ground truth terms as input. These methods require costly annotated training data from budzianowski-etal-2018-multiwoz. In contrast, our approach requires no training and conducts both term and relation extraction. We use vukovic-etal-2024-dialogue and finch-etal-2024-transforming as baselines.
@@ 2.2 Text-to-SQL
li2023llmDBinterface present a large benchmark for text-to-SQL queries on big databases and find LLMs underperform compared to humans. zhou_db-gpt_2024 apply LLMs to database tasks such as query rewriting and index tuning using automatic prompt generation, DB-specific pre-training, model design, and fine-tuning. yang-etal-2024-synthesizing enhance text-to-SQL parsing for smaller models by synthesising training data with larger models. allemang2024ontologies improve text-to-SPARQL Polleres2014sparql using a rule-based query checker that leverages the ontology, with an LLM repairing queries based on this feedback. stricker-paroubek-2024-shot use LLMs in a few-shot TOD setup by translating user queries into SQL to interact with tables in the underlying ontology. In contrast, we build a DB and ontology from scratch using text-to-SQL.
2.3 Dialogue Flow Induction
burdisso-etal-2024-dialog2flow propose Dialog2Flow (D2F) embeddings that map utterances to a latent space according to the actions they represent. These actions are relevant for dialogue flow and help unify action representations across datasets. agrawal-etal-2024-dialog induce dialogue flows unsupervisedly via clustering and merging, ensuring relevance, coverage, clarity, and non-redundancy. choubey-etal-2025-turning retrieve relevant conversations and generate dialogue workflows using a question–answer chain-of-thought prompt with an LLM, then evaluate them via dialogue simulation with workflow decomposition. These methods focus on inducing general system and user actions, closely related to intent and action prediction in TeQoDO. However, our approach also induces a fine-grained domain–slot–value hierarchy linking actions to underlying structured knowledge.
@@ 3.1 Problem Formulation
Given a TOD dataset $D=\{d_{1},\ldots,d_{k}\}$, each dialogue $d_{i}$ is comprised of alternating user and system turns $\{u_{1},s_{1},\ldots,u_{m_{i}},s_{m_{i}}\}$ containing information about user-queried entities. The goal is to induce an ontology $O_{D}$ for the dataset. The ontology is a directed graph $O_{D}=(V,E)$ with five node types $V=V_{\text{domain}}\cup V_{\text{user\_intent}}\cup V_{\text{system\_action}}\cup V_{\text{slot}}\cup V_{\text{value}}$, forming a three-level hierarchy. The edge set $E$ is a subset of edges between specific node types, $E\subseteq(V_{\mathrm{domain}}\times V_{\mathrm{slot}})\cup(V_{\mathrm{slot}}\times V_{\mathrm{value}})\cup(V_{\mathrm{user\_intent}}\times(V_{\mathrm{domain}}\times V_{\mathrm{slot}}))\cup(V_{\mathrm{system\_action}}\times(V_{\mathrm{domain}}\times V_{\mathrm{slot}})).$
@@ 3.2 Method
Our main prompting approach combines TOD modelling with SQL, as outlined below. We then detail the steps of our iterative ontology construction pipeline. We utilise an LLM that creates SQL queries, which are then executed.
@@
SQL is a query language for relational databases chamberlin1974sql. We use SQLite, [^2] https://www.sqlite.org a serverless relational DB management system suitable for small-scale databases and schema-based knowledge representation. SQL stores data in tables with columns and values, e.g., a “restaurant” table with a “price” column and “expensive” as a value. This structure aligns with TOD ontology formats, making it well-suited for ontology construction, as it reduces the prompt engineering needed for the LLM to induce the TOD ontology structure. For intents and actions, there are separate tables that store names of intent and action entries.
@@
We incorporate two concepts from modular task-oriented dialogue models (young2013pomdp) into our prompt to improve the quality of generated update queries. First, dialogue state tracking (DST) is used as a separate step to distinguish existing database content from new input, based on the current schema. Second, we prompt the model to consider dialogue success, guiding it to make updates that support achieving the user goal. For a more formal description, see Table˜8 in the appendix.
@@
See Figure˜3 for an overview of TeQoDO with example queries and results. The pipeline steps are given in Algorithm˜1 and are described in the following.
@@
To ensure structural and naming consistency in TeQoDO, the model is first prompted to query the current DB tables that are relevant for the current input dialogue $d_{i}$. This is done by giving the model the whole set of current tables in the DB as an additional input. Initially, as the ontology is built from scratch, the DB is empty, and thus no results are returned. The resulting table list $T=\{t_{1},\ldots,t_{n}\}$ is included with the dialogue. The model is prompted to query column info for tables relevant to the dialogue. We then append the column query results to the prompt so the model generates SELECT queries consistent with the DB state. Along with the results of the column information queries subsamples of explicit values a column contains are given to the model. In the final part of this step, it is prompted to generate SELECT queries to retrieve specific entity values from the dialogue. To reduce variation in column values (e.g. “Alexander Bed and Breakfast” vs “Alexander B & B”) and improve naming consistency in the DB, we also experiment with query results for similar tables, columns, and values. This decreases duplicate entries by increasing the chance of reusing existing values. In the final TeQoDO pipeline, we only utilise the column value subsamples.
@@
In this step, the model restructures DB query results into DST style labels. It predicts the current dialogue’s value for each table-column pair using the DB’s stored information. This improves the distinction between existing DB information and new information from the dialogue.
The model is prompted to summarise dialogue states using the DB’s table and column names. This summary includes information from both the user and the system. By predicting the state, the model condenses the dialogue’s information into a concise, model-determined format not fully detailed in the prompt. Notably, the model tracks the entire dialogue at once, not turn-by-turn. It must track all the mentioned information, not only what the user queries. Since the input is a whole dialogue, the model does not do traditional DST, but rather tracks the information that was mentioned in the dialogue in a format that is based on the current DB schema. The format here resembles DST labels with slot-value information.
Step 3: DB Update with Success
In the final step, the model generates database update queries based on the current DB state and dialogue information. It is prompted to create update queries considering DB query and DST results. Update queries include CREATE TABLE, ALTER TABLE to add columns, and UPDATE to add values. The model must follow the existing DB structure and generate consistent updates with existing tables, not create new ones. Additionally, the model is prompted to update the DB to support fulfilling the user’s goal by incorporating dialogue success into the prompt. Namely, this instruction is added to the update prompt: “so that the user’s goal expressed in the dialogue can be successfully fulfilled using only information stored in the database.
@@ 4.1 Datasets
The first dataset we employ is MultiWOZ 2.1 (eric-etal-2020-multiwoz), a large-scale, multi-domain dialogue dataset containing human-human conversations annotated with domain, slot, and value labels. It covers information such as hotel bookings, restaurant reservations, taxi services, and attractions. We utilise the test set with $1,000$ dialogues and $6$ domains.
@@ 7  Generate update queries $U_{d_{i}}$ for success using $p=p+ST_{d_{i}}$
Second, we use the schema-guided dialogue (SGD) dataset (rastogi2020towards). It is a diverse, multi-domain dataset for TOD modelling, featuring detailed schema annotations. It covers services like flight booking, calendar scheduling, banking, and media services, including unseen services at test time. We use the SGD test split for evaluation in the main results, containing $4,201$ dialogues and $18$ domains. We load these with ConvLab-3 zhu-etal-2023-convlab-local.
@@
To test whether TeQoDO applies beyond TOD data, we use the Wikipedia and arXiv datasets by lo2024ontologylearning. The Wikipedia test set contains $242,148$ titles and abstracts; its gold ontology has four hierarchy levels with $8,033$ nodes and $14,673$ edges. Topics include, for example, types of “injuries” or “legislative bodies”. Here, the input to TeQoDO are titles and summaries of Wikipedia articles rather than task-oriented dialogues. lo2024ontologylearning construct the dataset by running a breadth-first search from the root category “Main topic classifications” to depth 3, collecting titles and lead summaries of up to $5,000$ pages per category using the Wikipedia API [^3] https://en.wikipedia.org/w/api.php.
The arXiv test set has $27,630$ title–abstract pairs, two hierarchy levels with $61$ nodes and $61$ edges. Main topics are arXiv categories such as “Mathematics” with subcategories like “Commutative Algebra”, making this ontology more abstract and higher level than Wikipedia’s. Here, the input to TeQoDO are titles and abstracts of arXiv papers rather than TODs. The taxonomy is taken from the official site, and the corpus includes titles and abstracts of all arXiv papers from 2020–2022 with at least $10$ citations. These ontologies do not follow the domain–slot–value hierarchy, making SQL less intuitive.